myininaya (myininaya):

any ideas on how to evaluate \[\int_{0}^{1} \sqrt[7]{1-x^3}-\sqrt[3]{1-x^7} dx\]

myininaya (myininaya):

\[\int_{0}^{1} (\sqrt[7]{1-x^3}-\sqrt[3]{1-x^7} )dx\]

myininaya (myininaya):

its in the james stewart 6th edition calculus book the poblem plus section at the end of chapter 8 if anyone is interested

myininaya (myininaya):

and actually i wrote it backwards oh well

OpenStudy (anonymous):

stewart? never heard of it

myininaya (myininaya):

lol sorry for the delayed response my computer died

OpenStudy (anonymous):

in fact i have no idea how to do it. and i have only the stupid "essential" version with me

myininaya (myininaya):

we need the power of zarkon!

OpenStudy (jamesj):

Cogs turning ... I notice first that the two functions in the integrand are inverse of each other on [0,1]. But there's no magical theorem about the integral of (f -inv.f). None the less, context would be good--what is the author discussing immediately before this exercise? I also note that the function has one zero on (0,1); call it a. I've been trying to prove the integral from 0->a equals the negative of the integral a->1 without success yet.

myininaya (myininaya):

well its the problem plus section so anywhere from integration by parts, parital fractions, trig substitution, substitution there is also one thing the author used that i thought would be useful but had no luck in an example he proved \[\int\limits_{0}^{a}f(x)dx=\int\limits_{0}^{a}f(a-x) dx\]

OpenStudy (jamesj):

ok -- tv watching + problem solving, usually equals zero, but I'll give it a go.

myininaya (myininaya):

ok there is no hurry james i was thinking about giving it as a bonus assignment to my students but if i can't figure it out then i shouldn't make them try to figure it out lol i was just stumped by it i will also keep thinking on it

OpenStudy (amistre64):

\[\int_{0}^{1} (1-x^3)^{1/7}-{(1-x^7)^{1/3}} dx\] the xs get ya

OpenStudy (amistre64):

i wish i could remember, but there is something to do with a substitution with exponents

OpenStudy (amistre64):

you find a common denominator or some such and it goes z^(?) = some such thing

OpenStudy (amistre64):

\[z^7=1-x^3\] \[(1-z^7)^{1/3}=x\] \[7z^6 dz=-3x^2 du\]

OpenStudy (amistre64):

\[7z^6 dz=-3(1-z^7)^{2/3} du\] \[\frac{7z^6}{-3(1-z^7)^{2/3}} dz= dx\]

OpenStudy (amistre64):

...wish i could recall if this was the method im thinking of or if im just inventing it lol

myininaya (myininaya):

i don;t see anything wrong with it but i'm not sure if we can use it

OpenStudy (amistre64):

\[\int\frac{7z^6}{-3z(1-z^7)^{2/3}} dz\] \[\int\frac{-7z^6}{3z(1-z^7)^{2/3}} dz\] im not sure either :)

myininaya (myininaya):

still looks nasty

myininaya (myininaya):

this shouldn't be hard lol its just cal 2 cal 2>me :(

OpenStudy (amistre64):

7*3 = 21 so it was prolly something akin to z^{21} youd use or some such

myininaya (myininaya):

speak lagrange :)

OpenStudy (jamesj):

Ok. First, LEMMA Suppose f an increasing continuous function on [a,b]. Then \[\int\limits_{a}^{b} f^{-1} \ = \ b f^{-1}(b) - af^{-1}(a) - \int\limits_{f^{-1}(a)}^{f^{-1}(b)} f \ .\] PROOF Let \[P = \{ t_0, t_1, ..., t_n \}\] be a partition of [a,b] and let \[P' = \{ f^{-1}(t_0), ..., f^{-1}(t_n) \} .\] Then carefully writing out the sums with a diagram, we see that if L is a lower Riemann sum and U the upper sum, \[L(f^{-1}, P) + U(f, P') = \ b f^{-1}(b) - af^{-1}(a)\] The result now follows. Note also that we can obtain the same result with f decreasing on [a,b] by swapping L and U above. qed. Now, given all that, let \[f(x) = \sqrt[7]{1-x^3} \ \hbox{ and hence } \ f^{-1}(x) = \sqrt[3]{1-x^7}\] Consider \[I = \int\limits_{0}^{1} f^{-1} = 1.f^{-1}(1) - 0.f^{-1}(0) - \int\limits_{f^{-1}(0)}^{f^{-1}(1)} f\] \[ = 0 - 0 - \int\limits_{1}^{0} f = \int\limits_{0}^{1} f \] and therefore for our f \[\int\limits_{0}^{1} (f^{-1} - f\ ) = 0 \ .\] Hence \[\int\limits_{0}^{1} (\sqrt[3]{1-x^7} - \sqrt[7]{1-x^3} \ ) = 0 \ .\]

myininaya (myininaya):

wow james

OpenStudy (anonymous):

amazing!!!

OpenStudy (anonymous):

james is the man

myininaya (myininaya):

totally cute! too advance for cal 2 students though lol

OpenStudy (jamesj):

@Myininaya: where do you teach?

myininaya (myininaya):

its a small college in little rock Arkansas

OpenStudy (zarkon):

The result can also be shown by just using integration by parts for \[\int\limits_{a}^{b}f^{-1}(x)dx\] let \[u=f^{-1}(x)\text{ and }dv=dx\] and result follows fairly quickly.

OpenStudy (jamesj):

Really? I don't think it does, as du is messy.

OpenStudy (zarkon):

\[du=\frac{1}{f'(f^{-1}(x))}dx\text{ and }v=x\] \[\int\limits_{a}^{b}f^{-1}(x)dx=\left.xf^{-1}(x)\right|_{a}^{b}-\int\limits_{a}^{b}\frac{x}{f'(f^{-1}(x))}dx\] \[=bf^{-1}(b)-af^{-1}(a)-\int\limits_{a}^{b}\frac{x}{f'(f^{-1}(x))}dx\] \[u=f^{-1}(x) \Rightarrow f(u)=x\] since \[du=\frac{1}{f'(f^{-1}(x))}dx\] we have \[=bf^{-1}(b)-af^{-1}(a)-\int\limits_{u(a)}^{u(b)}f(u)du\] \[=bf^{-1}(b)-af^{-1}(a)-\int\limits_{f^{-1}(a)}^{f^{-1}(b)}f(u)du\] call x=u \[=bf^{-1}(b)-af^{-1}(a)-\int\limits_{f^{-1}(a)}^{f^{-1}(b)}f(x)dx\]

OpenStudy (jamesj):

Got it.

myininaya (myininaya):

james+zarkon=world domination by math